# Group-aware Parameter-efficient Updating for Content-Adaptive Neural Video Compression

### v1

@@ 4.2. Experimental Results
4.2.1. Compression Effectiveness on Videos.
In Fig. 4 and Table 1, we present quantitative comparisons for rate-distortion and BD-Rate (%) performance of our proposed method compared to existing video compression techniques across four video compression datasets. Our approach demonstrates superior performance over two contemporary content-adaptive NVC frameworks proposed by Tang et al. (Tang et al., 2023) and Lu et al. (Lu et al., 2020b), achieving bit-rate reductions of 59.01% and 65.57% on the average. Moreover, our experimental findings reveal a notable average bit-rate savings of 30.69% when compared to the conventional codec H.266/VVC, and a 2.51% improvement over the most recent NVC method, DCVC_DC.
@@
4.2.2. Compression adaptability on Medical Data.
In terms of MRI sequence compression, shown in Fig. 5 on the ACDC benchmark, our method exhibits a marked enhancement in compression performance. Compared to our baseline method, DCVC_DC, and the conventional video coding standard, H.266/VVC, our approach achieves bit-rate savings of 7.85% and 30.05%, respectively. This notable improvement underscores the adaptability of our method, highlighting its capacity to bridge the domain gap, which is a prevalent challenge in current NVC research, thereby allowing us to extend its applicability across varied domains.
@@
Table 1. BD-Rate (%) for PSNR on video benchmark. VTM-11.2 is used as the anchor. Negative values indicate bit-rate savings and positive values indicate additional bit-rate cost. The best and second-best performance are in bold and italic.
ClassB ClassC ClassD ClassE Avg
VTM-11.2 0.0 0.0 0.0 0.0 0.0
HM-16.20 42.86 19.51 28.42 58.76 37.39
DCVC 54.71 116.68 64.16 212.17 111.93
DCVC_DC -33.00 -8.42 -29.31 -39.03 -27.44
Lu et al. 106.54 107.61 102.56 267.96 146.17
Tang et al. 43.10 94.28 48.65 199.14 96.29
GPU (Ours) -34.88 -12.17 -32.13 -43.57 -30.68
@@ 4.3.1. Ablation I: Effectiveness of Adaptors
We undertook an ablation study utilizing the ACDC dataset on the machine with a single NVIDIA A100 GPU and AMD EPYC 7763 CPU. Here, we execute the online updating of all parameters associated with the encoder side (blue components in Fig. 3) within our content-adaptive NVC framework, designated as “Full Updating”. Fig. 7 illustrates the RD performance achieved by this full-updating strategy, while Table 2 details the associated computational complexity. Our findings reveal that, although the proposed PETL method incurs a negligible additional bit-rate cost (less than 2%) compared to the full updating approach, they significantly reduce the optimization requirement to less than 10% of the parameters. Moreover, these methods only necessitate 82.1% the optimizing time to train 1 iteration of 1 GoP (i.e., per epoch) required by the full updating strategy. Looking ahead, as more sophisticated techniques are integrated into NVC, it is anticipated that both the number of parameters and the encoding time could further increase, which means our architecture-agnostic updating methods could become increasingly practical.
Table 2. The computational complexity on ACDC of our method using encoding-side adapters to delta-tune the encoder (i.e., GPU) against the variant directly optimizing all modules within the encoding network (i.e., Full Updating).
Full Updating GPU (Ours)
#Optimizing Parameters (M) 2.91 0.23
Optimizing Time (s/Epoch) 2.85 2.34
4.3.2. Ablation II: Impact of frame number on online-updating efficiency
We also undertook a comparative analysis on the ACDC dataset to assess the impact of the number of frames updated during the online updating process. Specifically, our default configuration involves optimizing all frames (i.e., 32) within a GoP in one online updating session. This study extends our framework to include variations in the number of updated frames (i.e., 1, 10, and 20), each subset being smaller than the full GoP size. The results in Fig. 8 reveal that while the alternative strategies, which update only 10 and 20 frames, remain managing to improve upon our baseline method DCVC_DC, our default GPU updating the entire GoP outperforms these alternatives by achieving bit-rate savings of 7.98% and 3.76%, respectively. Notably, the variant only updating single frame experienced a considerable decline in performance, incurring an additional bit-rate cost of 21.86% over our GPU. This significant discrepancy underscores the fundamental differences between content-adaptive NIC and NVC, particularly highlighting the challenges associated with accumulating errors over time.
Figure 8. Performance of our methods against variants optimizing different numbers of frames during online updating.

### v2

@@ 4.2. Experimental Results
4.2.1. Video Compression.
In Fig. 4, we present quantitative comparisons for rate-distortion and BD-Rate (%) performance of our proposed method compared to existing video compression techniques across six video compression datasets. Our approach demonstrates superior performance over two contemporary content-adaptive NVC frameworks proposed by Tang et al.and Lu et al., achieving bit-rate reductions of 56.91% and 64.84% on the average. Moreover, our findings reveal a notable average bit-rate savings of 29.08% when compared to the conventional codec VTM, and a 2.25% improvement over the most recent method, DCVC_DC.
@@
4.2.2. Medical Volumetric Image Compression
In terms of MRI sequence compression, shown in Fig. 5 on the ACDC benchmark, our method exhibits a marked enhancement in compression performance. Compared to our baseline method, DCVC_DC, and the conventional video coding standard, VTM, our approach achieves bit-rate savings of 7.85% and 30.05%, respectively. This notable improvement underscores the adaptability of our updating method, highlighting its capacity to bridge the domain gap, which is a prevalent challenge in current NVC research, thereby allowing us to extend its applicability across varied domains.
@@ 4.3.1. Ablation I: Effectiveness of Adaptors
We undertook an ablation study utilizing the ACDC dataset. Here, we execute the online updating process of all parameters associated with the encoder side (blue components in Fig. 3) within our content-adaptive NVC framework, designated as “Full Updating”. Fig. 7 illustrates the RD performance achieved by this full-updating strategy, while Table 1 details the optimizing complexity of each coding component. Our findings reveal that, although the proposed PETL method incurs a negligible additional bit-rate cost (less than 2%) compared to the full updating approach, they significantly reduce the optimization requirement to less than 8% of the parameters. Looking ahead, as more sophisticated techniques are integrated into NVC, it is anticipated that both the number of parameters and the encoding time could further increase, which means our architecture-agnostic updating methods could become increasingly practical.
Table 1. The number of optimizing parameters (in Millions) for our method (i.e., GPU) against the variant directly optimizing all encoding-side parameters (i.e., Full Updating).
Encoding Components Full Updating GPU (Ours)
Motion Estimation 0.96 $4\times 10^{-4}$
Motion Encoder 0.43 0.05
Motion Hyper-prior 0.18 $2\times 10^{-3}$
Contexture Encoder 0.90 0.17
Contexture Hyper-prior 0.44 $9\times 10^{-3}$
Overall 2.91 0.23
4.3.2. Ablation II: Impact of P-frame number on online-updating efficiency
We also undertook a comparative analysis on the ACDC dataset to assess the impact of the number of P-frames updated during the online updating process. Specifically, our default configuration involves optimizing all P-frames (i.e., 31) within a GoP in one online updating session. This study extends our framework to include variations in the number of updated P-frames (i.e., 1, 10, and 20), each subset being smaller than the full GoP size. The results in Fig. 8 reveal that while the alternative strategies, which update only 10 and 20 P-frames, remain managing to improve upon our baseline method DCVC_DC, our default GPU updating the entire GoP outperforms these alternatives by achieving bit-rate savings of 7.98% and 3.76%, respectively. Notably, the variant only updating single P-frame experienced a considerable decline in performance, incurring an additional bit-rate cost of 21.86% over our GPU. This significant discrepancy underscores the fundamental differences between content-adaptive NIC and NVC, particularly highlighting the challenges associated with accumulating errors over time.
Figure 8. Performance of our methods against variants optimizing various numbers of P-frames during online updating.
@@ 5. Conclusion
Acknowledgements.
This work was supported in part by Startup funds from The University of Newcastle, Australia, in part by the Hong Kong Research Grants Council General Research Fund (17203023), in part by The Hong Kong Jockey Club Charities Trust under Grant 2022-0174, in part by the Startup Fund and the Seed Fund for Basic Research for New Staff from The University of Hong Kong, and in part by the funding from UBTECH Robotics.